### v1

Abstract
The advent of foundation models have revolutionized various fields, enabling unprecedented task accuracy and flexibility in computational linguistics, computer vision and other domains. Attention mechanism has become an essential component of foundation models, due to their superb capability of capturing correlations in a sequence. However, attention results in quadratic complexity in memory and compute as the context length grows. Although many fusion-based exact attention acceleration algorithms have been developed for datacenter-grade GPUs and accelerators leveraging multi-core parallelism and data locality, yet it remains a significant challenge to accelerate attention on resource-constrained edge neural accelerators with limited compute units and stringent on-chip caches. In this paper, we propose a scheme for exact attention inference acceleration on memory-constrained edge accelerators, by parallelizing the utilization of heterogeneous compute units, i.e., vector processing units and matrix processing units. Our method involves scheduling workloads onto these different compute units in a multi-tiered tiling scheme to process tiled vector workloads and matrix workloads in attention as two streams, respecting the workload dependencies. We search for tiling factors to maximize the parallelization of both compute units while considering I/O overhead, and propose a proactive cache overwrite strategy to avoid undesirable cache spills in reality. Extensive results based on open-sourced simulation frameworks show up to 2.75$\times$ speedup and 54% reduction in energy consumption as compared to the state-of-the-art attention fusion method (FLAT) in the edge computing scenario. Further experiments on a real-world edge neural processing unit demonstrate speedup of up to 1.76$\times$ for attention as compared to FLAT, without affecting model output accuracy.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Foundation models Vaswani et al. (2017); Kitaev et al. (2020); Kaplan et al. (2020); Peebles & Xie (2023); Li et al. (2024a) have driven recent advancements in generative AI on edge devices such as smartphones, especially in AI agents Zhang et al. (2023); Wang et al. (2024); Fan et al. (2025), large language models (LLMs) Radford et al. (2018); Ouyang et al. (2022); Glaese et al. (2022); Mehta et al. (2024) and text-to-image diffusion models Poole et al. (2022); Esser et al. (2024). Central to these models is the attention mechanism, which captures long-range dependencies between tokens, but incurs quadratic memory and computational complexity due to pairwise token interactions. Deploying these models is challenging, especially on resource-constrained edge devices with limited on-chip cache and processing power.
Significant efforts have been made to accelerate attention computation through software fusion techniques on datacenter-grade hardware. For cloud servers, multi-core parallelism Shoeybi et al. (2019); Rasley et al. (2020); Narayanan et al. (2021); Kwon et al. (2023); Liu et al. (2023); Cho et al. (2024) and efficient utilization of on-chip SRAM in GPUs Kirk et al. (2007) are employed to enhance performance. FlashAttention Dao et al. (2022); Dao (2023); Shah et al. (2024); dao ; Hong et al. (2023) related methods design I/O-aware exact attention speedup algorithms, leveraging GPU CUDA cores and on-chip SRAM to minimize access to High Bandwidth Memory (HBM), saving memory and reducing runtime. FuseMax Nayak et al. (2024) uses Einsums and a spatial array accelerator, employing ping-pong scheduling to overlap MatMul and softmax operations. While FlashAttention-3 and FuseMax parallelize MatMul and softmax on multi-core architectures, these cloud-based acceleration methods cannot be applied on resource-constrained edge accelerators directly, which have a limited number of processing units and on-chip memory.
To speed up attention inference on edge devices, current methods mainly leverage graph fusion Ivanov et al. (2021); Niu et al. (2021); Aminabadi et al. (2022); Mei et al. (2023) to restrict or reduce data transfers between off-chip and on-chip memory. TVM Chen et al. (2018a) utilizes an automated schedule optimizer to improve execution for a given neural network. Although TVM’s auto-scheduler is designed for general purposes, the limitation is that it does not fuse MatMul and softmax operators in the attention block. oneDNN Li et al. (2024b) tackles the fusion of MatMul and softmax with graph fusion templates and microkernels to accelerate attention on Intel CPUs, while FLAT Kao et al. (2023) uses row-granularity tiling and on-chip cache to alleviate the bandwidth bottleneck to access off-chip memory, achieving speedup and energy savings. Similarly, proprietary technologies like NVIDIA TensorRT Nvi and Apple CoreML app (b; c; a) claim to leverage graph fusion for attention acceleration.
Although these advancements promote fusion and data locality on edge devices, to the best of our knowledge, most existing works in the public domain (e.g., FLAT) still execute the workloads including matrix multiplication (MatMul) and softmax sequentially, which achieves suboptimal latency. It remains a significant challenge to execute heterogeneous workloads, including MatMul workloads that typically run on the multiplier accumulator (MAC) compute unit and softmax workloads that typically rely on the vector (VEC) unit, in parallel on edge accelerators with limited cores and thus limited or no multi-core parallelism opportunities. Furthermore, the limited on-chip memory demands careful memory management schemes to prevent cache overflow and redundant computation.
In this paper, we introduce Memory Aware Stream Processing Attention (MAS-Attention) to accelerate attention computation on resource-constrained edge devices. MAS-Attention employs a semi-synchronous parallelization strategy to simultaneously utilize the heterogeneous MAC compute unit and vector compute unit on a neural accelerator in a pipelined parallel fashion, minimizing bubbles and optimizing the cachce management to improve attention inference efficiency. Our contributions can be summarized as follows:
  • We propose a novel stream processing scheme that parallelizes tiled MatMul and softmax workloads through a semi-synchronous pipelining process. Prior works only parallelize the computing and I/O processes while still executing operators sequentially. In contrast, we aim to schedule all operators in the attention mechanism onto the heterogeneous computing units, to achieve parallel execution by scheduling the stream of MatMul workloads on the MAC unit and stream of softmax workloads on the VEC unit while satisfying data dependencies between tiled workloads.
  • We employ a multi-tiered tiling scheme for MAS-Attention dataflow, that accommodates key hardware constraints and software parameters. This scheme employs fine-grained sub-matrix tiling for MatMul and row-granularity tiling for softmax operations. Using search strategies, we identify optimal tensor tiling factors to balance workloads efficiently within the stream processing scheme through offline auto-tuning across different attention workloads and hardware configurations.
  • A proactive buffer overwrite strategy is further introduced to maintain efficiency with limited on-chip buffer capacity, especially for longer input sequences. This approach selectively overwrites specific MAC unit data to prioritize softmax completion with fine-grained control, minimizing data reloading. It ensures data dependencies, maintains operand integrity, and prevents pipeline stalls or reverting to prior rounds.
We extensively evaluate MAS-Attention across attention layers in transformer-based models, including different variants of BERT Devlin et al. (2018), Llama3-8B Touvron et al. (2023), T5 Raffel et al. (2020), ViT Dosovitskiy et al. (2020) and XLM Lample & Conneau (2019). For simulations, we utilize a modified TileFlow Zheng et al. (2023) to define the edge spatial accelerator architecture, software mapping, and search space exploration, while Timeloop Parashar et al. (2019) and Accelergy Wu et al. (2019) are used to estimate latency and energy consumption. Additionally, we test MAS-Attention on real hardware, using a Huawei MatePad Pro 13.2 with a DaVinci Liao et al. (2019) NPU. On the simulated edge device, MAS-Attention achieves up to $2.75\texttimes$ speedup and $54\%$ reduction in energy consumption compared to the state-of-the-art FLAT algorithm. Similar improvements in speedup and energy savings are also observed on the actual edge NPU hardware, further validating MAS-Attention’s effectiveness.
2 Related Work
Sequential Attention Execution: The Layer-Wise attention computation processes operations sequentially. This method relies on transferring intermediate results between off-chip and on-chip memory, creating a memory-bound workflow that poses significant deployment challenges on edge devices with limited memory bandwidth.
Approximate Attention Acceleration Methods: For approximate acceleration methods of transformer-based foundation models, methods like palletization Cho et al. (2021); Tabani et al. (2021); Wang et al. (2020a); app (b), quantization Liu et al. (2021); Lin et al. (2021); Wang et al. (2022); Li et al. (2022); Piao et al. (2022); Yao et al. (2022); Li & Gu (2023); Yu et al. (2023), pruning Mao et al. (2021); Peng et al. (2021); Yu et al. (2022b; a), and knowledge distillation Sun et al. (2019); Wang et al. (2020c; b); Ganesh et al. (2021); Huang et al. (2024); Gupta et al. (2024) compress model size by reducing parameters or transferring knowledge from larger models, achieving memory efficiency and faster inference.
Exact Attention Acceleration Methods: In cloud environments, exact acceleration methods Dao et al. (2022); Dao (2023); Shah et al. (2024); dao ; Hong et al. (2023); Patel et al. (2023); Nayak et al. (2024) leverage parallel computation on multi-core architectures to speed up attention mechanism. For instance, FlashAttention and FlashAttention-2 optimize dataflow for attention computation on NVIDIA A100 GPUs by dividing query, key and value inputs into smaller tiles and loading them from high-bandwidth memory (HBM) to on-chip SRAM, reducing data movement for large intermediate outputs and exploiting GPU CUDA core parallelism. FlashAttention-3 further enhances parallelism using ping-pong scheduling to overlap MatMul and softmax operations within warp groups on NVIDIA H100 GPUs. In contrast, FuseMax uses Einsums to design a spatial array accelerator for fused attention calculation in cloud environments, overlapping MatMul and softmax computations and improving spatial PE array utilization.
Due to limited computing cores, resource-constrained edge devices rely on graph-fusion-based kernels Gao et al. (1993); Kjolstad et al. (2017); Chen et al. (2018b); Baghdadi et al. (2019); goo ; Zhou & Yang (2022)—such as oneDNN Li et al. (2024b) for CPUs and FLAT Kao et al. (2023)—to accelerate attention computations by fusing operators and retaining intermediate results on-chip, which reduces DRAM and off-chip memory access overhead. FLAT employs a row-based attention fusion strategy for TPUs Jouppi et al. (2017; 2020) and spatial accelerators Kwon et al. (2018); Chen et al. (2019), including edge devices. By loading rows of query into on-chip memory, FLAT performs corresponding MatMul and softmax row-wise computations on-chip and writes the output rows directly to off-chip memory, thus mitigating memory-bound limitations by minimizing large data transfers. However, previous attention acceleration methods overlook the heterogeneous computing characteristics between MatMul and softmax, which run on MAC and VEC units, missing an opportunity for parallelization that could further reduce latency and energy consumption.
3 MAS-Attention Overview
While prior exact attention acceleration methods for resource-constrained edge devices, such as oneDNN Li et al. (2024b) and FLAT Kao et al. (2023), enhance data locality and reduce memory access overhead through operator fusion, they still execute tiled MatMul and softmax operators sequentially, missing the chance for parallel execution within the attention mechanism. In our work, we leverage the heterogeneous computing capabilities of edge devices to achieve parallel execution of tiled MatMul and softmax for the exact attention acceleration. Our method further minimizes the latency with this parallelization scheme while reducing I/O and redundant memory access with a novel multi-tiered tiling scheme and a proactive memory-aware buffer management, making it advantageous even for single inference requests in AI scenarios on resource-constrained edge devices.
Heterogeneous Workloads of Attention Mechanism: On resource-constrained edge devices, heterogeneous computing is often used to perform computation and memory access concurrently. However, prior edge-based attention acceleration methods have not explored the heterogeneous nature of MatMul and softmax workloads within the attention mechanism. Given their distinct computational characteristics, the compute-intensive MatMul operation runs on the MAC unit, while the element-wise softmax operation is processed on the VEC unit. Leveraging this heterogeneity, enables parallel execution of MatMul and softmax computations, providing further acceleration of the attention mechanism.
Hardware-Software Co-design Scheduling on Resource-Constrained Edge Devices: Given the limited computing cores and on-chip memory, efficiently scheduling tiled MatMul and softmax operators with parallel execution in the attention workload requires consideration of both hardware parameters (e.g., L1 and L0 memory sizes, MAC and VEC core counts) and software parameters (e.g., MatMul and softmax workload shapes). To address this challenging hardware-software co-design scheduling problem, we propose a novel multi-tiered tiling scheme that accommodates both short and long sequence lengths while enhancing the utilization of on-chip processing units. Specifically, we introduce sub-matrix tiling granularity for MatMul and row-wise tiling granularity for softmax workloads. This approach creates distinct tiling search spaces for different workloads, allowing for higher search efficiency. We employ advanced search algorithms like MCTS to conduct offline searches for obtaining optimal tiling parameters across various attention workloads and hardware configurations. Our tiling scheme and search algorithm aim to balance MAC and VEC operations in a fused, pipelined, semi-synchronous attention computation, maximizing processing unit utilization, minimizing idle time, and reducing I/O and redundant memory access to ultimately optimize inference latency and energy consumption.
Memory-aware Optimizations for the Limited Shared On-chip Memory: While our multi-tiered tiling scheme allocates search budgets for different workloads within the attention mechanism and enhances the efficiency of the search algorithm, limited search budgets can lead to locally optimal tiling parameters, particularly for long input sequences with extensive search spaces. Additionally, the constrained shared on-chip memory in edge devices complicates the scheduling of parallelized MatMul and softmax workloads. To address the potential for sub-optimal tiling parameters and better utilize limited on-chip memory, we introduce an innovative proactive buffer overwrite strategy. This memory-aware optimization features guardian mechanisms that proactively overwrite selected on-chip buffered data, balancing data refetching and redundant computation against cache overflow. It prioritizes critical operators for timely completion while ensuring correct data dependencies within the pipelined dataflow.
Figure 1: Dataflow comparison between FLAT and MAS-Attention: FLAT executes tiled stages sequentially, while MAS-Attention performs MatMul and softmax operations semi-synchronously in parallel, maximizing compute utilization and significantly enhancing overall performance.
4 Methodology
Given the query, key and value matrices, $Q,K,V\in\mathbb{R}^{{B}\times{H}\times{N}\times{E}}$, where $B$ is the batch size, $H$ is the number of heads, $N$ is the sequence length and $E$ is the embedding size, the attention output $O$ is computed through the following steps:
$C=Q{K^{T}}\in\mathbb{R}^{{B}\times{H}\times{N}\times{N}},$ (1)
$P=\mbox{softmax}(C)\in\mathbb{R}^{{B}\times{H}\times{N}\times{N}},$ (2)
$O=PV\in\mathbb{R}^{{B}\times{H}\times{N}\times{E}},$ (3)
where softmax is applied to every row of $QK^{T}$.
To efficiently perform these computations on resource-limited spatial accelerators, we propose a semi-synchronous MAC-VEC parallel execution scheme. Our method is applicable to a wide range of spatial accelerators that have at least one MAC unit for matrix multiplications and one VEC unit for element-wise operations. Our scheme is achieved through the strategic scheduling and pipelining of two MatMul operations alongside a single Softmax operation, as illustrated in Figure 1. This approach allows the three operators to concurrently process different tiles within the same computation round, thereby accelerating the attention mechanism. Additionally, we leverage advanced heuristic search algorithms to optimize the tiling sizes across all memory levels within our dataflow. These algorithms adaptively tune the tiling parameters based on input dimensions, workload characteristics, and pipelining criteria to ensure a balanced distribution of workloads across compute units. We also implement an on-chip memory management strategy that selectively overwrites non-essential data to free up memory resources, prioritizing Softmax computation for longer sequences while ensuring the subsequent recovery of interrupted MatMul operations. Detailed descriptions of these strategies are provided in the following.
4.1 Stream Processing Mechanism
We propose a stream processing scheme to handle continuous streams of tiled MatMul and Softmax workloads. There are two streams of tiled tasks: one for tiled MatMul computation (defined in Algorithms 2 and 4) and another for tiled Softmax computation defined in Algorithm 3. These streams are scheduled in a pipelined fashion to overlap tiled MatMul-Softmax computations, as illustrated in Figure 1.
Our approach operates at a row granularity, where the input matrix $Q$ is divided into smaller chunks along the batch, head, and sequence dimensions, resulting in row-wise sub-matrices denoted as $Q_{i}$. This granularity is driven by the inherently row-wise nature of the Softmax operation, aligning the processing scheme with Softmax’s requirements. Iterations thus proceed based on the segmented sequence dimension of the query, allowing for efficient parallelism. The detailed stream processing scheme is outlined in Algorithm 1, where there are warm-up, regular, and finalize computation rounds.
Algorithm 1 MAS-Attention
1:  Require: $\mathbf{Q,K,V}\in\mathbb{R}^{B\times H\times N\times E}$ in DRAM; Parameters $B_{b},H_{h},N_{Q},N_{K,V}\in\mathbb{R}$
2:  Divide $\mathbf{Q}$ into $T_{r}=\left\lceil\frac{B}{B_{b}}\right\rceil\times\left\lceil\frac{H}{H_{h}}\right\rceil\times\left\lceil\frac{N}{N_{Q}}\right\rceil$ blocks $\mathbf{Q_{1},...,Q_{r}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times E}$
3:  Divide $\mathbf{O}$ into $T_{r}=\left\lceil\frac{B}{B_{b}}\right\rceil\times\left\lceil\frac{H}{H_{h}}\right\rceil\times\left\lceil\frac{N}{N_{Q}}\right\rceil$ blocks $\mathbf{O_{1},...,O_{r}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times E}$
4:  Allocate $(B,H,N,E)$ for $\mathbf{O}$ in DRAM
5:  Call Alg. 2: $\mathbf{C_{1}}\leftarrow\mathbf{Q_{1}K^{T}}$
6:  $i\leftarrow 2$
7:  while $i\leq T_{r}$ do
8:   if $i=2$ then
9:    Parallel Execution:
10:     Call Alg. 2: $\mathbf{C_{2}}\leftarrow\mathbf{Q_{2}K^{T}}$
11:     Call Alg. 3: $\mathbf{P_{1}}\leftarrow\text{Softmax}(\mathbf{C_{1}})$
12:   else
13:    Parallel Execution:
14:     Call Alg. 4: $\mathbf{O_{i-2}}\leftarrow\mathbf{P_{i-2}V}$
15:     Call Alg. 3: $\mathbf{P_{i-1}}\leftarrow\text{Softmax}(\mathbf{C_{i-1}})$
16:     Wait for completion of Alg. 4 then:
17:      Call Alg. 2: $\mathbf{C_{i}}\leftarrow\mathbf{Q_{i}K^{T}}$
18:   end if
19:   $i\leftarrow i+1$
20:  end while
21:  Finalize:
22:   Parallel Execution:
23:    Call Alg. 4: $\mathbf{O_{i-2}}\leftarrow\mathbf{P_{i-2}V}$
24:    Call Alg. 3: $\mathbf{P_{i-1}}\leftarrow\text{Softmax}(\mathbf{C_{i-1}})$
25:   Wait for completion of Alg. 3 then:
26:    Call Alg. 4: $\mathbf{O_{i-1}}\leftarrow\mathbf{P_{i-1}V}$
27:  return $\mathbf{O}$
In the warm-up computation round, we use the MAC unit to compute the first tile for the first MatMul operator as $C_{1}=Q_{1}K^{T}$. Then, we use the VEC unit to compute the first tile for the Softmax operator as $P_{1}=\mbox{Softmax}(C_{1})$ and use the MAC unit to compute the second tile for the first MatMul operator as $C_{2}=Q_{2}K^{T}$ in parallel. Then we enter the regular computation rounds, as shown in lines 13-17 of Algorithm 1. For iterations $i\geq 3$, the MAC unit computes the tile for the final MatMul operator as $O_{i-2}=P_{i-2}V$. Meanwhile, the VEC unit computes the tile for the Softmax operator as $P_{i-1}=\mbox{Softmax}(C_{i-1})$. While the tiled Softmax task is being processed, the MAC unit computes the tile for the first MatMul operator as $C_{i}=Q_{i}K^{T}$ upon completion of $O_{i-2}$. Lastly, in the finalize computation round, the MAC unit computes the last tile for the final MatMul operator as $O_{i-1}=P_{i-1}V$ after the VEC unit computes the last tile of the Softmax operator as $P_{i-1}=\mbox{Softmax}(C_{i-1})$.
Our pipelined attention mechanism operates in a semi-synchronous manner. During a regular computation round, there is no data dependency among workloads, allowing the two tiled MatMuls and Softmax to be executed in parallel by the MAC and VEC units, respectively. However, within each computation round, data dependencies must be carefully managed to ensure the correctness of the computation. This semi-synchronous MAC-VEC parallelism for MatMul-Softmax computations significantly reduces the latency of the attention mechanism.
4.2 MAS-Attention Tiling Scheme
We introduce a multi-tiered tiling strategy for MAS-Attention dataflow. For matrices $K$, $P$ and $V$, used in the MatMul operations in Equation 1 and Equation 3, a fine-grained sub-matrix tiling is applied. This approach is crucial, especially when the sequence length is significantly longer than the embedding dimension ($N\gg E$), as it helps address the constraints of limited on-chip memory. Without such tiling, handling the matrix $K$ in $C_{i}=Q_{i}K^{T}$ and the matrices $P_{i}$ and $V$ in $O_{i}=P_{i}V$ becomes problematic due to excessive memory demands. For intermediate tensors $C_{i}$ and $P_{i}$ used in the Softmax operation in Equation 2, a row-granularity tiling is employed, aligning with the inherent row-wise nature of Softmax to maintain computational correctness.
We establish a comprehensive search space for tiling parameters across the memory hierarchy of the targeted hardware, focusing on dimensions such as batch size ($B$), number of attention heads ($H$), query sequence length ($N_{Q}$), and key/value sequence lengths ($N_{K,V}$). The search for optimal tiling parameters is influenced by three key factors: the detailed workload of attention mechanism, the specific scheduling of MAS-Attention, and the input size. These parameters are defined at each memory level to ensure efficient off-chip and on-chip memory operations while considering the interaction between computation and memory usage. This approach aims to identify optimal or near-optimal tiling configurations that maintain computational efficiency throughout the stream processing of MAS-Attention. To effectively navigate this search space, we use Genetic Algorithms and Monte Carlo Tree Search (MCTS) for the simulated edge device, and Grid Search for the edge device with a DaVinci DNN Accelerator.
The Genetic Algorithm evolves a population of tiling configurations through selection, crossover, and mutation, iteratively refining the configurations to optimize execution cycles and power consumption. MCTS, on the other hand, builds a search tree where each node represents a potential tiling configuration, balancing exploration of new configurations with exploitation of high-performing ones to converge toward optimal or near-optimal solutions. On the DaVinci DNN Accelerator, Grid Search systematically evaluates all possible configurations, leveraging its compatibility with the hardware’s structured memory model. These algorithms iteratively assess various tiling configurations, simulating different tile shapes and sizes to determine those that optimize execution cycles and minimize power consumption.
After retrieving the optimal tiling parameters from the search, Algorithm 2 performs the tiled MatMul computation of $C_{i}$, using sub-matrices $Q_{i}$ and finer-grained sub-tiles of $K$. The algorithm reads blocks of $Q_{i}$ and sub-blocks of $K$ from DRAM to on-chip memory, where the MatMul operation $Q_{i}K^{T}$ is executed to generate $C_{i}$. The resulting $C_{i}$ is retained on-chip for subsequent operations.
Algorithm 2 Produce $C_{i}\leftarrow Q_{i}K^{T}$
1:  Require: $\mathbf{Q_{i}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times E}$, $\mathbf{K}\in\mathbb{R}^{B\times H\times N\times E}$ in DRAM; $B_{b},H_{h},N_{Q},N_{K,V}\in\mathbb{R}$; $i=[b_{b}:b_{e},h_{b}:h_{e},n_{b}:n_{e}]$
2:  Select $i^{th}$ set of batch and head from $\mathbf{K}$ as $\mathbf{K_{i}}=\mathbf{K}[i[0],i[1],:,:]\in\mathbb{R}^{B_{b}\times H_{h}\times N\times E}$
3:  Divide $\mathbf{K_{i}}$ into $T_{c}=\left\lceil\frac{N}{N_{K,V}}\right\rceil$ blocks $\mathbf{K_{i,1}},\ldots,\mathbf{K_{i,c}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{K,V}\times E}$
4:  Allocate $(B_{b},H_{h},N_{Q},N)$ for $\mathbf{C_{i}}$, $(B_{b},H_{h},N_{Q},E)$ for $\mathbf{Q_{i}}$, and $(B_{b},H_{h},N_{K,V},E)$ for $\mathbf{K_{i,j}}$ in on-chip memory
5:  Load $\mathbf{Q_{i}}$ from DRAM to on-chip memory
6:  for $1\leq j\leq T_{c}$ do
7:     Load $\mathbf{K_{i,j}}$ from DRAM to on-chip memory
8:     On-chip compute $\mathbf{C_{i,j}}=\mathbf{Q_{i}K_{i,j}^{T}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times N_{K,V}}$
9:     Write $\mathbf{C_{i,j}}$ to on-chip memory as $j^{th}$ block of $\mathbf{C_{i}}$
10:  end for
Algorithm 3 handles the tiled softmax computation. It processes the on-chip $C_{i}$ matrix by dividing it into smaller row-wise blocks, aligned with the row-wise nature of the softmax operation. Each block undergoes the softmax steps: identifying the maximum value, subtracting it, exponentiating, summing, and normalizing to produce $P_{i}$. The $P_{i}$ blocks are kept on-chip to ensure efficient data access for final Matmul computation.
Algorithm 3 Produce $P_{i}\leftarrow C_{i}$
1:  Require: $\mathbf{C_{i}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times N}$ in on-chip memory; $B_{b},H_{h},N_{Q}\in\mathbb{R}$; $i=[b_{b}:b_{e},h_{b}:h_{e},n_{b}:n_{e}]$
2:  Divide $\mathbf{C_{i}}$ into $T_{l}=N_{Q}$ blocks $\mathbf{C_{i,1}},\ldots,\mathbf{C_{i,l}}\in\mathbb{R}^{B_{b}\times H_{h}\times 1\times N}$
3:  Allocate $(B_{b},H_{h},N_{Q},N)$ for $\mathbf{P_{i}}$ in on-chip memory
4:  for $1\leq j\leq T_{l}$ do
5:     On-chip compute $P_{i,j}=\text{Softmax}(\mathbf{C_{i,j}})\in\mathbb{R}^{B_{b}\times H_{h}\times 1\times N}$
6:     Write $\mathbf{P_{i,j}}$ to on-chip memory as $j^{th}$ block of $\mathbf{P_{i}}$
7:  end for
Algorithm 4 handles the tiled MatMul computation of $O_{i}$. Both $P_{i}$ and $V_{i}$ are divided into finer-grained blocks to manage large sequence lengths. In each iteration, a block of $V_{i}$ is loaded from DRAM to on-chip memory, while a corresponding block of $P_{i}$ is already available on-chip. The block-wise multiplication $P_{i,j}V_{i,j}$ is performed iteratively, accumulating results into $O_{i}$. Once all iterations are complete, $O_{i}$ is written back to off-chip memory.
Algorithm 4 Produce $O_{i}\leftarrow P_{i}V$
1:  Require: $\mathbf{P_{i}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times N}$ in on-chip memory; $\mathbf{V}\in\mathbb{R}^{B\times H\times N\times E}$, $\mathbf{O}\in\mathbb{R}^{B\times H\times N\times E}$ in DRAM; $B_{b},H_{h},N_{Q},N_{K,V}\in\mathbb{R}$; $i=[b_{b}:b_{e},h_{b}:h_{e},n_{b}:n_{e}]$
2:  Select $i^{th}$ set batch and head from $\mathbf{V}$ as $\mathbf{V_{i}}=\mathbf{V}[i[0],i[1],:,:]\in\mathbb{R}^{B_{b}\times H_{h}\times N\times E}$
3:  Divide $\mathbf{V_{i}}$ into $T_{c}=\left\lceil\frac{N}{N_{K,V}}\right\rceil$ blocks $\mathbf{V_{i,1}},\ldots,\mathbf{V_{i,c}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{K,V}\times E}$
4:  Divide $\mathbf{P_{i}}$ into $T_{c}=\left\lceil\frac{N}{N_{K,V}}\right\rceil$ blocks $\mathbf{P_{i,1}},\ldots,\mathbf{P_{i,c}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times N_{K,V}}$
5:  Allocate $(B_{b},H_{h},N_{Q},E)$ for $\mathbf{O_{i}}$, and $(B_{b},H_{h},N_{KV},E)$ for $\mathbf{V_{i,j}}$ in on-chip memory
6:  On-chip initialize $\mathbf{O_{i}}=(0)_{B_{b}\times H_{h}\times N_{Q}\times E}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times E}$
7:  for $1\leq j\leq T_{c}$ do
8:     Load $\mathbf{V_{i,j}}$ from DRAM to on-chip memory
9:     On-chip compute $\mathbf{O_{i}}=\mathbf{O_{i}}+\mathbf{P_{i,j}V_{i,j}}\in\mathbb{R}^{B_{b}\times H_{h}\times N_{Q}\times E}$
10:  end for
11:  Write $\mathbf{O_{i}}$ to off-chip memory as $i^{th}$ block of $\mathbf{O}$
4.3 Proactive Overwrite Strategy for Optimized Memory Utilization
The tiling parameters obtained from heuristic search algorithms, such as Genetic Algorithm and Monte Carlo Tree Search, may not always yield optimal results. Due to the complexity of the search space and the heuristic nature of these algorithms, there is a possibility of suboptimal configurations, which can impact the efficiency and correctness of stream processing. To mitigate these potential inefficiencies and ensure robust performance across a variety of workloads and scenarios, we introduce a selective overwrite strategy. This proactive approach enables the system to adaptively manage on-chip memory by selectively overwriting specific non-essential data when memory constraints arise.
During the computation of $P_{i}$, if the on-chip memory reaches capacity, impeding further calculations, two cases may arise. First, as shown in Figure 2, if the MAC unit is engaged in processing $P_{i-1}V$, $P_{i}$ will overwrite the $V$ matrix on chip and stop the MAC from continuing its operation, resulting in no more writes from the MAC unit to on-chip buffer. Second, as shown in Figure 3, if the MAC unit is occupied with $Q_{i+1}K^{T}$, $P_{i}$ will overwrite the $K$ matrix on chip, thereby interrupting the MAC unit’s process and preventing any further writes to the on-chip memory. Once the final result of $P_{i}$ is fully calculated and stored on chip, the MAC unit can resume its process by reloading either the $V$ or $K$ matrix from DRAM to on-chip memory if it was overwritten and redoing the MatMul calculation.
Figure 2: Selective Overwriting of $V$ Matrix to Halt MatMul Operation in MAS-Attention’s Memory Strategy.
Figure 3: Selective Overwriting of $K$ Matrix to Halt MatMul Operation in MAS-Attention’s Memory Strategy.
The rationale is that maintaining the integrity of critical operands is essential to the efficiency of the pipeline. Preserving and finishing $P_{i}=\mbox{softmax}(C_{i})$ is crucial as the softmax operation stores its results only on chip and depends on $C_{i}=Q_{i}K^{T}$ which was obtained from on-chip memory, hence overwriting $P_{i}$ cannot be remedied by reloading it from DRAM. In contrast, this is not the case for $K$ and $V$ matrices, overwriting of which can be remedied by reloading the over-written tensors from DRAM without stalling the pipeline computation rounds.
This strategy ensures efficient use of on-chip memory and computational resources, and our careful data overwrite method makes the impact of the increased number of DRAM reads on our overall latency and energy savings unnoticeable. By carefully managing memory overwrites and reloading only essential data, we strike a balance between maximizing parallelism and maintaining computational efficiency, ultimately leading to improved performance and energy efficiency.
5 Experiments
5.1 Experimental Setup
This section provides details on the simulation and modeling tools utilized, describes the hardware specifications, and outlines the experimental workloads and baseline algorithms used for analysis. We conduct a comprehensive evaluation of the proposed method, comparing it against state-of-the-art attention fusion and acceleration techniques tailored for spatial accelerators in edge environments. This includes an assessment of performance on foundation model workloads suited for edge deployment.
Simulation and Modeling Tools: To simulate our experiments, we employed Timeloop Parashar et al. (2019) and Accelergy Wu et al. (2019) to measure the latency and energy consumption, also we modified TileFlow Zheng et al. (2023) to define the edge spatial accelerator, software mapping for attention inference, and search space exploration. The custom edge hardware architecture designed for simulation operates at a frequency of 3.75GHz and features 16nm technology, two cores each containing a MAC and a VEC unit, and a hierarchical memory system as depicted in Figure 4. The designed DRAM has a bandwidth of 30GB/sec and a total size of 6GB. The L1 cache has connection to DRAM and L0 register file and has a storage of 5MB . The Processing Elements (PEs) in MAC and VEC units, organized in 16x16 and 256 mesh respectively, have access to L0 register file. These parameters for the hardware architecture were determined after various stress tests of the hardware. Our simulations were conducted on a system equipped with an Intel(R) Xeon(R) Gold 6140 CPU @ 2.30GHz, utilizing a single thread in execution. Additionally, we evaluated our algorithm on a real hardware, specifically Huawei MatePad Pro 13.2 with DaVinci DNN Accelerator, to validate its practical applicability and performance.
Figure 4: Simulated Edge Architecture Design
Workloads: The workload for our experiments focuses on the inference of attention layers in various transformer-based networks, including different variants of BERT Devlin et al. (2018), Llama Touvron et al. (2023), T5 Raffel et al. (2020), ViT Dosovitskiy et al. (2020), and XLM Lample & Conneau (2019), as detailed in Table 1. We selected a diverse set of networks with varying attention layer dimensions to ensure a comprehensive evaluation. Additionally, each workload in our experiments undergoes a rigorous golden data check for all methods, including our proposed approach, ensuring that all methods pass this validation.
Layer-Wise: This approach represents the unfused baseline for attention inference. In this method, $C=QK^{T}$ is fully computed first, followed by the the softmax function on the entire matrix $C$ to yield $P=\text{Softmax}(C)$. Once $P$ is fully computed, the final output of the attention unit, $O=PV$, is then calculated. All these operations occur sequentially and without fusion.
Soft-Pipe: For comparison, we also design a baseline algorithm that only pipelines the first MatMul and the softmax. It divides $Q$ and $K$ into smaller chunks, fuses and pipelines MatMul of $C=QK^{T}$ with $\text{Softmax}(C)$. In each iteration, rows of $Q$ ($Q_{i}$) are loaded into on-chip memory to compute the corresponding rows of $C$, where $C_{i}=Q_{i}K^{T}$. Then, the corresponding rows of $P$ are computed on-chip with $P_{i}=\text{Softmax}(C_{i})$. While $P_{i}$ is being calculated, $C_{i+1}$ can be computed simultaneously. The resulting $P$ values are stored back to DRAM, and once the computation of $P$ is complete, the final output of the attention unit, $O=PV$, is calculated sequentially.
FLAT: $Q$, $K$, and $V$ matrices are divided into smaller chunks, and all attention operations are fused on-chip and computed sequentially. In each iteration, rows of $Q$ ($Q_{i}$) are loaded into on-chip memory to compute the corresponding rows of $C$, where $C_{i}=Q_{i}K^{T}$. Then, the corresponding rows of $P$ are computed on-chip with $P_{i}=\text{Softmax}(C_{i})$. Finally, the corresponding rows of $O$, where $O_{i}=P_{i}V$, are computed on-chip and written back to off-chip memory.
TileFlow: In this approach, $Q$, $K$, and $V$ are divided into smaller chunks, and all operations in the attention unit are fused and pipelined Zheng et al. (2023). However, since Zheng et al. (2023) does not provide further implementation details, we implemented the algorithm to the best of our knowledge based on the available information. Specifically, we replicated TileFlow’s tiling and pipelining approach by dividing matrices into sub-tiles that fit within on-chip memory, fusing MatMul and softmax operations with pipeline execution. This implementation approximates TileFlow’s behavior, ensuring our evaluation aligns with its intended operational characteristics.
Table 1: Network Configuration and Hyper-Parameters
Network Name #Heads #Seq Hidden size EmbK,V
BERT-Base & T5-Base 12 512 768 64
BERT-Large & T5-Large 16 512 1024 64
BERT-Small 8 512 512 64
Llama3-8B & T5-3B (T5-XL) 32 512 4096 128
T5-Mini & T5-Small 8 512 256 32
ViT-B/14 12 196 768 64
ViT-L/14 16 196 1024 64
ViT-H/14 16 196 1280 80
ViT-B/16 12 256 768 64
ViT-L/16 16 256 1024 64
ViT-H/16 16 256 1280 80
XLM 8 512 1024 128
5.2 Execution Time Analysis
5.2.1 Analysis on Simulated Hardware
Table 2 presents a detailed analysis of execution cycles and speedup ratios for MAS-Attention compared to other methods across all tested networks. The data highlights that MAS-Attention consistently achieves superior performance, with speedup factors ranging from 1.02×1.02$\times$1.02 × to 8.50×8.50$\times$8.50 × over Layer-Wise, and up to 5.25×5.25$\times$5.25 × over Soft-Pipe, 2.75×2.75$\times$2.75 × over FLAT, and 2.00×2.$0$ over TileFlow methods. The geometric mean of these speedup values—5.09×5.09$\times$5.09 ×, 2.78×2.78$\times$2.78 ×, 1.70×1.70$\times$1.70 ×, and 1.31×1.31$\times$1.31 × respectively—demonstrates MAS-Attention’s overall efficiency in reducing execution time. This substantial performance improvement underscores MAS-Attention’s effectiveness as an advanced solution for optimizing computational efficiency in attention mechanisms.
Table 2: Cycles and Speedup Comparisons Across Networks for Different Methods
Network Name Cycles ($10^{6}$) Speedup (MAS-Attention vs. Others)
Layer-Wise Soft-Pipe FLAT TileFlow MAS-Attention Layer-Wise Soft-Pipe FLAT TileFlow
BERT-Base & T5-Base 3.637 2.064 1.573 0.799 0.786 4.63 2.63 2.00 1.02
BERT-Large & T5-Large 5.505 2.753 1.835 1.311 1.049 5.25 2.63 1.75 1.25
BERT-Small 2.753 1.376 0.918 0.655 0.524 5.25 2.63 1.75 1.25
Llama3-8B & T5-3B (T5-XL) 12.845 8.389 4.719 5.243 4.194 3.06 2.00 1.13 1.25
T5-Mini & T5-Small 2.228 1.180 0.721 0.328 0.262 8.50 4.50 2.75 1.25
ViT-B/14 0.612 0.381 0.266 0.263 0.151 4.06 2.53 1.77 1.75
ViT-L/14 1.242 0.508 0.354 0.351 0.201 6.19 2.53 1.77 1.75
ViT-H/14 1.355 0.558 0.405 0.439 0.251 5.40 2.23 1.61 1.75
ViT-B/16 1.081 0.590 0.426 0.249 0.197 5.50 3.00 2.17 1.27
ViT-L/16 1.311 0.786 0.524 0.332 0.262 5.00 3.00 2.00 1.27
ViT-H/16 1.376 0.852 0.590 0.414 0.328 4.20 2.60 1.80 1.26
XLM 4.194 2.097 1.180 1.311 1.049 4.00 2.00 1.13 1.25
Geometric Mean - - - - - 5.09x 2.78x 1.70x 1.31x
5.2.2 Analysis on Real Hardware
Figure 5 shows the analysis of normalized execution time for Layer-Wise, Soft-Pipe, FLAT, and MAS-Attention methods on Huawei MatePad Pro 13.2 with DaVinci DNN Accelerator. MAS-Attention achieves substantial performance improvements, with speedups ranging from 1.94×1.94$\times$1.94 × to 3.50×3.50$\times$3.50 × over Layer-Wise, 1.35×1.35$\times$1.35 × to 2.87×2.87$\times$2.87 × over Soft-Pipe, and 1.30×1.30$\times$1.30 × to 1.76×1.76$\times$1.76 × over FLAT. The geometric mean speedups are 2.33×2.$3$, 1.73×1.73$\times$1.73 ×, and 1.42×1.42$\times$1.42 ×, respectively. It is worth noting that TileFlow was not included in this analysis as its implementation details were not fully described in Zheng et al. (2023), which limited us from deploying it on this edge device. Overall, the data validates MAS-Attention’s effectiveness in enhancing computational efficiency on real hardware.
Figure 5: Normalized Execution Time Comparison Across Networks for Different Methods on Huawei MatePad Pro 13.2 with DaVinci DNN Accelerator
5.3 Power and Energy Analysis
Table 3 presents a comprehensive analysis of energy consumption and savings achieved by MAS-Attention compared to other methods across various networks. The data reveals that MAS-Attention consistently demonstrates significant energy efficiency improvements, with savings ranging from $1.98\%$ to $75.00\%$ compared to Layer-Wise, $36.83\%$ to $66.08\%$ compared to Soft-Pipe, $0.02\%$ to $54.03\%$ compared to FLAT, and $36.83\%$ to $65.05$ compared to TileFlow. The geometric mean of these savings—$52.97\%$, $63.07\%$, $18.55\%$, and $53.16\%$ respectively—highlights MAS-Attention’s overall effectiveness in reducing energy consumption. This substantial reduction underscores MAS-Attention’s potential for optimizing energy efficiency in resource limited environments.
Table 3: Energy Consumption and Savings Comparisons Across Networks for Different Methods.
Network Name Energy Consumption ($10^{9}$ nJ) Energy Savings (MAS-Attention vs. Others)
Layer-Wise Soft-Pipe FLAT TileFlow MAS-Attention Layer-Wise Soft-Pipe FLAT TileFlow
BERT-base & T5-Base 37.208 49.607 12.656 27.598 12.405 66.67% 75.00% 1.98% 55.05%
BERT-large & T5-Large 28.105 65.672 21.112 38.065 16.944 39.69% 74.20% 19.75% 55.49%
BERT-small 20.218 24.336 10.556 19.032 8.359 58.65% 65.64% 20.80% 56.08%
Llama3-8B & T5-3B (T5-XL) 179.309 186.463 63.252 147.502 63.241 64.73% 66.08% 0.02% 57.12%
T5-Mini & T5-Small 12.434 11.269 8.744 7.512 4.746 61.83% 57.90% 45.71% 36.83%
ViT-B/14 3.720 7.376 2.803 4.136 1.903 48.87% 74.21% 32.11% 54.00%
ViT-L/14 5.539 7.335 5.648 7.428 2.596 53.13% 64.61% 54.03% 65.05%
ViT-H/14 6.585 9.120 4.741 6.783 3.162 51.98% 65.34% 33.27% 53.38%
ViT-B/16 5.323 5.828 3.350 7.119 3.239 39.16% 44.42% 3.34% 54.49%
ViT-L/16 9.403 6.984 6.316 9.402 4.218 55.14% 39.61% 33.21% 55.14%
ViT-H/16 11.160 15.414 6.803 11.475 5.156 53.81% 66.55% 24.22% 55.09%
XLM-Base 35.786 46.485 15.813 36.876 15.584 56.45% 66.47% 1.45% 57.74%
Geometric Mean - - - - - 52.97% 63.07% 18.55% 53.16%
Note: Based on some literature studies, “nJ” (nanojoules) is used as the unit for energy consumption reported by Accelergy.
In addition, we provide an energy consumption breakdown for each network on all algorithms as shown in Figure 6, focusing on Off-Chip (DRAM) and On-Chip (L1, L0) memories, and PEs in MAC and Vector units.
5.3.1 Off-Chip Memory Energy Consumption
Compared to Layer-Wise and Soft-Pipe methods, MAS-Attention significantly reduces off-chip energy consumption by minimizing DRAM accesses and eliminating the need to store intermediate $C$ and $P$ matrices off-chip. However, MAS-Attention’s off-chip energy consumption in some cases is slightly higher than FLAT due to the need of reloading $K$ and $V$ matrices in the case of them being overwritten by the selective overwriting mechanism during pipelining. Soft-Pipe consumes more energy than MAS-Attention as it stores the $P$ matrix back to DRAM, but less than Layer-Wise as it does not store the $C$ matrix to DRAM.
5.3.2 On-Chip Memory Energy Consumption
Layer-Wise, Soft-Pipe and TileFlow usually consumes much more on-chip energy compared to MAS-Attention, indicating less efficient on-chip memory utilization. FLAT also show higher energy consumption than MAS-Attention but generally lower than Layer-Wise, Soft-Pipe and TileFlow.
5.3.3 PEs Energy Consumption
Energy consumption in PEs remains constant across different algorithms for each network, as the actual computation required by different algorithms is the same, with differences only in the scheduling process.
Figure 6: Energy Consumption Breakdown for DDR, L1, L0 memories and PEs within MAC and VEC units Across Networks using Different Methods
5.4 DRAM Access Analysis
Since the FLAT method is most comparable to MAS-Attention in terms of both cycle and energy performance, we will focus on comparing the DRAM access between these two algorithms.
5.4.1 DRAM Write Operations
Both MAS-Attention and FLAT algorithms exhibit an identical number of write operations to DRAM. This uniformity arises because both algorithms confine their DRAM write operations to the final result of the attention block ($O$), eschewing the need to write intermediate results to DRAM. Instead, these intermediate results are processed entirely on-chip, thereby minimizing off-chip memory accesses and enhancing overall efficiency.
5.4.2 DRAM Read Operations
Across the tested workloads, MAS-Attention matches FLAT in DRAM read operations but surpasses it for specific networks. Notably, for BERT-Base & T5-Base (1.5×1.5$\times$1.5 ×), BERT-Large & T5-Large (1.5×1.5$\times$1.5 ×), and Llama3-8B & T5-3B (1.49×1.49$\times$1.49 ×), MAS-Attention shows increased DRAM read operations. This phenomenon arises because MAS-Attention requires reloading specific data chunks, particularly $K$ and $V$ matrices, which may have been overwritten during the pipelining stages on-chip. These matrices are reloaded from DRAM to resume the halted MAC operations, allowing the attention mechanism to maintain data dependencies and continue processing seamlessly. While this incurs additional DRAM reads, the proactive buffer overwriting mechanism maintains efficient on-chip memory usage and pipelined execution integrity, with total cycle counts and energy consumption still outperforming all other baselines.
5.5 Limitations
On the simulated edge hardware, MAS-Attention can handle a maximum sequence length of approximately 1 million tokens in half precision (FP16), which is half the maximum sequence length that FLAT can handle. The computation of $P_{i}$ happens in parallel with either $O_{i-1}$ or $C_{i+1}$. Since Softmax operates row-wise, at least one row is used in the computation of $P_{i}$. In the case of $P_{i}$ computed in parallel with $O_{i-1}=P_{i-1}V$, $O_{i-1}$ requires at least one entire row of $P_{i-1}$ to be calculated. Also, in the case of $P_{i}$ computed in parallel with $C_{i+1}=Q_{i+1}K^{T}$, one entire row of $C_{i+1}$ is computed and written on-chip. In both scenarios, on-chip memory should have the capacity for either $P_{i}$ and $P_{i-1}$ or $P_{i}$ and $C_{i+1}$. In the case of half precision with a sequence length of 1M, one row of $P_{i}$, $P_{i-1}$, and $C_{i+1}$ consumes 2MB each on-chip, which fits within the 5MB on-chip cache size in either scenario. Since FLAT does not employ such a pipelining scheme and operates sequentially, it can handle a sequence length of 2 million tokens. In this condition, one row of $P_{i}$ consumes 4MB on-chip, which can be managed by the 5MB on-chip cache size in the simulated edge device.
6 Conclusion & Future Work
In this paper, we propose MAS-Attention dataflow to accelerate attention mechanism on resource-constrained edge devices. Our approach uses a stream processing scheme to execute tiled MatMul and Softmax workloads in a pipelined manner, with MAC and VEC units operating in parallel. A multi-tiered tiling strategy ensures balanced workloads for efficient pipelined attention execution. Additionally, our proactive buffer overwrite strategy enhances on-chip memory utilization by freeing up buffer space when it runs out of memory, such as with longer input sequences. While this strategy increases off-chip memory reads, MAS-Attention achieves superior speedup and energy savings over previous methods like Layer-wise, Soft-Pipe, FLAT, and TileFlow, on both simulated and real edge devices.
Future work will extend MAS-Attention to support training, which adds complexity in backpropagation that challenges efficient workload management on resource-constrained edge devices.